[figure: 9.7 cm]

REENTRANCE IN TRICRITICAL FIELD BEHAVIOUR OF DILUTED ISING METAMAGNET

M. Žukovič and T. Idogaki

Department of Applied Science, Faculty of Engineering, Kyushu University, Japan

Critical behaviour of diluted anisotropic antiferromagnets in a field is still a subject of discussion. Previous both experimental and theoretical works have suggested that there is a tricritical point (TCP), separating the temperature-field (T − H) phase space into a second-order (T > T_{TCP} and H < H_{TCP}) and first-order (T < T_{TCP} and H > H_{TCP}) transition regions. The TCP sustains until dilution reaches a certain value p_{t}, which has been argued to be equal or higher than the two-dimensional percolation threshold p_{c}(d = 2). At this concentration, beyond which no first-order transition can exist, the temperature T_{TCP} goes to zero while the field H_{TCP} remains finite. We have performed Monte Carlo calculations for the quenched site-diluted spin-\frac{1}{2} Ising metamagnet with alternating ferromagnetic planes coupled antiferromagnetically. Our previous effective field calculations suggested that there might be a certain field region where the transition curve at a fixed value of the field crosses over from the second-order region to the firstorder one and then again back to the second-order one (the reentrant phenomenon). Motivated by these calculations, our aim was to check the possibility of such a behaviour. We found indeed evidence of the first-order transition at concentration as low as p = 0.6 at a field intensity considerably lower than the tricritical field intensity H_{TCP}(p_{t}) = p_{t}z_{2}|J_{2}| (z_{2} is the number of nearest neighbours coupled with the central spin by J_{2} < 0). This result strongly suggests that H_{TCP} − p curve is a double-valued function of H, i.e. a possibility of the reentrant phenomenon.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M.Žukovič

Address for correspondence : Department of Applied Science, Faculty of Engineering, Kyushu University Hakozaki 6-10-1, Higashi-ku, 812-8581 Fukuoka, Japan

Email address : milantap@mbox.nc.kyushu-u.ac.jp